A 12-V battery, a 12-ohm resistor and a 4-ohm resistor are connected as

shown. The current in the 12-ohm resistor is ____ that in the 4-ohm resistor.


1
Expert's answer
2022-04-04T09:05:07-0400

Answer

Now think these resistance are connected in series then

Total resistance

R=12+4=16"\Omega"

So current is same in both resistance so

"I =\frac{V}{R}\\=\frac{12}{16}\\=0.75A"